OpenStudy (anonymous):

i have a test really soon in ratios,standard form,and proportions, can anyone explain them to me? :/

OpenStudy (anonymous):

ask doubts pl.

OpenStudy (anonymous):

Ratio- a:b, a/b Standard form- ax^2+bx+c Proportion- a/b=c/d

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

A ratio is really a fraction

OpenStudy (anonymous):

there are 565 sheets of paper in a book. how many sheets of paper are there in 200 of these books? give your answer in standard form

OpenStudy (anonymous):

200x565

OpenStudy (anonymous):

200*565= answer

OpenStudy (anonymous):

Sorry I have no calculator :P

OpenStudy (anonymous):

113000

OpenStudy (anonymous):

standard form?

OpenStudy (anonymous):

the meaning of that is?

OpenStudy (anonymous):

how do we write 113000 in standard form?

OpenStudy (anonymous):

Yes that means not in scientic notation

OpenStudy (anonymous):

oh ok :S im so blonde sometimes

OpenStudy (anonymous):

and theres part 2

OpenStudy (anonymous):

a pile of 565 sheets of paper is 25 mm high. calculate the thickness of 1 sheet of paper in standard form

OpenStudy (anonymous):

565/25

OpenStudy (anonymous):

are you sure you are givin me the correct answers cz it seems rly simple